I am trying to solve for N where I know (439.2/N)/(643.6/N) = 0.6824. How can I get N?
Answer by checkley77(12844)

(439.2/N)/(643.6/N)=.6824
(439.2/N)*(N/643.6)=.6824 THE N VALUES CANCEL OUT.
439.2/643.6=.6824
.6824=.6824